# ADA-Track: End-to-End Multi-Camera 3D Multi-Object Tracking with Alternating Detection and Association

### v1

@@ 4 Experiments
4.1 Experiment Setup
@@
DETR3D TBA-Baseline [62] TBA 0.321 1.448 0.452 0.283 474 15269 43828 57595
TBD-Baseline TBD 0.350 1.427 0.467 0.308 944 14316 42980 57973
MUTR3D [62] TBA 0.294 1.498 0.427 0.267 3822
DQTrack [34] TBD 0.367 1.351 1120
ADA-Track (ours) ours 0.378 1.391 0.507 0.343 981 15443 38466 62450
STAR-Track [18] TBA 0.379 1.358 0.501 0.360 372
ADA-Track-long (ours) ours 0.392 1.375 0.523 0.363 897 14711 36945 64055
PETR TBA-Baseline [62] TBA 0.407 1.357 0.511 0.369 271 14401 39487 62139
TBD-Baseline TBD 0.452 1.275 0.540 0.405 1035 13873 35269 65593
PF-Track [46] (w/o extension) TBA 0.453 642
PF-Track [46] TBA 0.479 1.227 0.590 0.435 181
ADA-Track (ours) ours 0.479 1.246 0.602 0.430 767 15385 31402 69728
Table 1: Results on the nuScenes validation set. We group methods by using the DETR3D or PETR detector. Tracking-by-attention and tracking-by-detection baselines are shown in the first two columns of each group. The remaining part compares our method (shown in light gray background) with existing works. denotes total training epochs of 48. indicates annotations from future frames are required.
@@
ADA-Track (ours) 0.456 1.237 0.559 0.406 834
Table 2: Comparison with state-of-the-art end-to-end multi-camera 3D MOT approaches on the nuScenes test split. The bottom part compares our method with other query-based methods that use DETR3D or PETR. The top part shows other methods.
Dataset
@@
Implementation details
To compare with most multi-camera MOT works [62, 18, 46, 34], we validate our approach using two query-based detectors: DETR3D [55] and PETR [40]. For DETR3D experiments, input images are full-resolution $1600\times 900$ and we use ResNet-101 [21] as the backbone with an FPN [36]. For PETR experiments, we crop the images to $1600\times 640$ and use VoVNetV2 [29] with FPN [36] as image feature extractor. We initialize the model using the corresponding single-frame detector checkpoints pre-trained for 24 epochs. We then train our tracker for 24 epochs on sampled mini-sequences which contain $T=3$ frames. For all DETR3D experiments, we freeze the weights of the image backbone and FPN following STAR-Track [18]. For PETR, we only freeze the image backbone following PF-Track [46]. All models are trained using a cosine-annealing schedule with an initial learning rate of $2e^{-4}$ and an AdamW [41] optimizer with a weight decay of $1e^{-2}$. All DETR3D experiments are trained on four V100 GPUs, while all PETR experiments use eight A100 GPUs. Each GPU holds one batch element.
@@ 4.2 Paradigm Comparison
Results
Table 1 shows the comparison on the nuScenes validation split, where the baselines are shown in the top part of each detector group. For both detectors, the TBD-Baselines achieve significantly higher AMOTA than the TBA-Baselines. On top, ADA-Track outperforms the TBD-Baselines by 2.8%P AMOTA for DETR3D and 2.7%P for PETR, respectively, while achieving a considerably higher recall. Considering the secondary metrics, we observe that the TBA-Baseline produces fewer IDS than the TBD-Baseline and ADA-Track, i.e. both methods with explicit association. However, TBA-Baseline achieves a low IDS by tracking only the easy cases, resulting in a substantially higher FN and lower MOTA. Compared to TBD-Baseline, ADA-Track again improves MOTA by a much higher TP and lower FN. These observations show the importance of the decoupled queries with explicit association as in the TBD-Baseline and in ADA-Track, which produces distinguishable query representations for both tasks and thus improves their performances. ADA-Track further improves over the TBD-Baseline by fully utilizing the task inter-dependency using alternating detection and association.
@@ 4.3 Comparison with Existing Works
We compare ADA-Track with existing works based on TBA or TBD in the remaining part of each detector group in Table 1. Due to an implementation issue, MUTR3D [62] reported a lower performance than our TBA-Baseline, which is a reproduction of MUTR3D with fewer training epochs and fixed backbone during training. DQ-Track [34] also uses decoupled queries and a sophisticated learned association module following TBD. ADA-Track outperforms it by 1.2%P AMOTA, which again highlights the effectiveness of our alternating detection and association design. The training of STAR-Track [18] requires an initialization on a pre-trained MUTR3D checkpoint, which results in a total training epoch of 48. For a fair comparison, we additionally train our model for 48 epochs, denoted as ADA-Track-long, which outperforms STAR-Track by 1.3%P AMOTA.
Comparing based on the PETR detector, we achieve on-par performance (0.479 AMOTA) with PF-Track [46]. However, PF-Track [46] is a joint tracking and prediction method, utilizing a track extension module to replace low-confidence detections with predicted trajectories when outputting tracking results, which additionally requires supervision from future frames. Our ADA-Track is a pure tracking method but still achieves the same AMOTA. Compared to PF-Track without track extension, ADA-Track achieves considerably higher AMOTA by 2.6%P.
We compare ADA-Track with end-to-end methods on the test split in Table 2, where we train ADA-Track based on DETR3D [55] with VoVNetV2-99 [29] backbone on both training and validation set. Compared to query-based methods using DETR3D or PETR in the second part of Table 2, ADA-Track achieves 0.456 AMOTA and 1.237 AMOTP, outperforming the recent state-of-the-art methods STAR-Track [18] and PF-Track [46] by 1.7%P and 2.2%P AMOTA, respectively. Compared to other non-query-based methods, ADA-Track also achieves the best performance, improving CC-3DT [20] that used a stronger BEVFormer [35] detector by 4.6% AMOTA.
In addition, we highlight that ADA-Track can be combined with many components proposed in existing tracking-by-attention works. For instance, the Past and Future Reasoning in [46] or the Latent Motion Model in [18] improve the query embedding or the motion update during query propagation, while we use the simple approach as in MUTR3D. These extensions can be seamlessly integrated into our framework and further improvements are expected.
4.4 Ablation Study
Number of training frames
Table 3 shows the impact of varying the training sample length $T$. We observe that an increasing sample length increases the AMOTA and many secondary metrics. A particularly substantial improvement occurs when increasing from a sample length of 2 to 3 frames, resulting in a notable 3.2%P increase in AMOTA. The reason is that the autoregressive training scheme becomes active when $T\geq 3$, allowing association results to propagate into subsequent frames. This enables optimization through gradients across multiple frames to optimize the whole sequence, significantly boosting robustness during inference. A further increase in training frames from 3 to 4 yields marginal improvements but a considerable computational and memory overhead during training. Thus, we use $T=3$ as the default setting.
Joint optimization
One of the main contributions of this work is the introduction of an alternating detection and association paradigm, where both tasks iteratively inform each other layer-by-layer for joint optimization. To assess its effectiveness, we combine the predictions of the bounding boxes and the association scores from different decoder layers. The first part of Table 4 illustrates using bounding boxes from the last layer alongside association scores from decoder layers 1 to 5. A notable increase in AMOTA is observed when using association scores from the second instead of the first layer. Using association scores from higher layers leads to gradually increased AMOTA. We also observe a considerable reduction in IDS. Together, this shows the iterative improvements of the association. The second part of Table 4 presents results using boxes from various layers combined with association scores from the last layer. We can observe a similar tendency in AMOTA increase with higher layers as in the first part, where the most significant increase occurs from the first to the second layer. In addition, using box predictions in later layers results in a notable improvement in AMOTP, confirming the iterative optimization of localization precision. Overall, combining outputs from distinct layers yields gradual performance increases when outputs from higher layers are used. This observation shows that the relevance between box and association predictions is not only constrained within the same layer but across different layers, which validates the iterative optimization through stacked layers of both tasks.
@@ 4.4 Ablation Study
2 0.346 1.397 0.477 0.295 1099
3 0.378 1.391 0.507 0.343 981
4 0.379 1.405 0.516 0.345 816
Table 3: Ablation study on the training sample length $T$.
@@
6 1 0.342 1.420 0.485 0.298 2063
6 2 0.369 1.399 0.485 0.330 1392
6 3 0.372 1.399 0.498 0.341 1088
6 4 0.373 1.398 0.508 0.338 1090
6 5 0.374 1.397 0.501 0.343 968
1 6 0.279 1.465 0.398 0.248 1431
2 6 0.364 1.413 0.480 0.329 1197
3 6 0.368 1.406 0.489 0.337 1023
4 6 0.373 1.399 0.513 0.342 1023
5 6 0.373 1.390 0.501 0.339 924
6 6 0.378 1.391 0.507 0.343 981
Table 4: Ablation study on combining bounding box and association outputs from different decoder layers.
Feature update weight
We analyze the impact of the feature update as discussed in Section 3.2. Here we generalize the feature update to a weighted average, i.e. $q_{\text{T},i}=w_{\text{T}}q_{\text{T},i}^{(L_{d})}+(1-w_{\text{T}})q_{\text{D},j}^{(L_{d})}$, where $w_{t}$ is the update rate. As shown in Table 5, substituting track query features with detection query features ($w_{t}=0$) yields the optimal performance, surpassing other values by 0.7% to 1.5% in AMOTA. This can be attributed to the fact that detection queries inherently incorporate track query features within their representation through edge-augmented cross-attention. Thus, an additional merging between detection and track query features is unnecessary.
@@
0 0.378 1.391 0.507 0.343 981
0.3 0.367 1.405 0.474 0.330 865
0.5 0.371 1.393 0.508 0.322 904
0.7 0.363 1.418 0.501 0.325 967
1.0 0.365 1.398 0.500 0.325 928
Table 5: Ablation study on the feature update weight $w_{t}$.
@@
A 0.365 1.390 0.489
B 0.343 1.423 0.489
C 0.378 1.391 0.507
Table 6: Ablation study on box update. use track box denotes that boxes predicted from track queries are used as output. track as query denotes that track queries are regarded as queries and detection queries as keys in the edge-augmented cross-attention.
Box update
Both track and detection queries predict bounding boxes, resulting in two candidates for each instance when two queries are associated. We use the box from the detection side to represent this associated pair and validate this choice in Table 6. As shown in the first row (Exp. A), selecting the track box results in an AMOTA decrease of 1.3%P compared to our approach (Exp. C), which shows the better quality of detection boxes. One could argue that detection queries are further refined in the edge-augmented cross-attention but track queries are not. In the second row of Table 6, we additionally reverse the edge-augmented cross-attention by using tracks as queries and detections as keys. This leads to an additional AMOTA decrease of 2.2%P. This finding also verifies the necessity of the decoupled queries for decoupled tasks, where using detection queries to locate objects and associate them to track queries is more effective than relying on track queries to locate objects.
5 Conclusion
We presented a novel query-based multi-camera 3D multi-object tracking approach, termed ADA-Track. We observed that decoupling the detection and association task while simultaneously leveraging the synergies between these tasks is the key to accomplish high-quality tracking. In line with this finding, we proposed a paradigm that conducts both, detection and association, in an alternating manner. In addition, we proposed a learned association module based on edge-augmented cross-attention which can be seamlessly integrated into any query-based decoder. Extensive experiments show the effectiveness of our approach compared to other paradigms, while achieving state-of-the-art performance on the nuScenes tracking benchmark.
Acknowledgement
@@
Supplementary Material
Appendix A Model Details
We provide the details of the model architectures as well as some design choices for the experiments with the DETR3D [55] and PETR [40] detector.
DETR3D
For all DETR3D-based [55] experiments, we use ResNet-101 [21] as the backbone. A FPN [36] image neck is attached to the ResNet-101 which outputs multi-scale feature maps $\{C_{2},C_{3},C_{4},C_{5}\}$ with downsampling rates $\{\frac{1}{4},\frac{1}{8},\frac{1}{16},\frac{1}{32}\}$ for the detection head. The detection head consists of 6 transformer decoder layers. Each decoder layer has a query-to-query self-attention, a DETR3D-based query-to-image cross-attention, and an edge-augmented cross-attention. The box regression and classification heads with two-layer MLPs are attached to the output of the DETR3D-based query-to-image cross-attention. All decoder layers have an embedding dimension of 256 and a feed-forward dimension of 512. Following DETR3D [55], the query embeddings and query position encodings of the detection queries are randomly initialized, and the initial reference points are estimated from their initial position encoding using a linear layer.
PETR
For all PETR-based [40] experiments, we utilize VoVNetV2-99 [29] as backbone and a FPN [36] image neck. The FPN feature map $C_{5}$ is upsampled and fused with $C_{4}$, producing the final single-scale feature map with the downsampling rate $\frac{1}{16}$ for the detection head. The architecture of the detection head is the same as in the DETR3D-based experiments, except that the query-to-image cross-attention is based on PETR [40] with 3D position encodings. The embedding dimension is 256 and the feed-forward dimension is 2048. In contrast to [55], PETR [40] generates query position encodings using the uniformly initialized reference points while initializing detection queries. In the query propagation phase, we adhere to this setting to update the query position encodings using the updated reference point positions at each timestamp. We found this design choice to be essential for ensuring the effectiveness of the PETR-based ADA-Track.
Appendix B Additional Experiments
We provide additional experiments of ADA-Track to validate our system design. All experiments are evaluated on the nuScenes validation set with the DETR3D detector.
bicycle bus car motorcycle pedestrian trailer truck average
TBA-Baseline 0.549 0.553 0.706 0.499 0.597 0.276 0.532 0.530
ADA-Track 0.551 0.675 0.725 0.545 0.613 0.336 0.571 0.574
Table A1: IDF1 for all categories on the nuScenes validation set.
Params FLOPS Inference time
TBA-Baseline 59.73 M 1024 G 296 ms
TBD-Baseline 63.72 M 1025 G 297 ms
ADA-Track 63.73 M 1054 G 308 ms
Table A2: Complexity and runtime analysis based on DETR3D with ResNet-101 backbone and image resolution of $1600\times 900$. Runtime is measured on an RTX 2080ti.
B.1 IDF1 Analysis

### v2

@@ 5 Experiments
5.1 Implementation Details
We evaluate our approach using three image backbones, ResNet-101 [65], VoVNet-99 [66] and ViT-Large [67], and two query-based detectors, DETR3D [7] and PETR [8]. The detection head of all experiments consists of 6 transformer decoder layers. The model is initialized from the corresponding single-frame detector checkpoints pre-trained for 24 epochs. We then train the end-to-end tracker for another 24 epochs on sampled mini-sequences that contain $T=3$ frames. All models are trained using a cosine-annealing schedule with an initial learning rate of $2e^{-4}$ and an AdamW [68] optimizer with a weight decay of $1e^{-2}$. All experiments with ResNet-101 backbone are conducted on four V100 GPUs, while other backbones use eight A100 GPUs. Each GPU holds one batch element. All ablation studies are conducted using DETR3D with ResNet-101.
For all DETR3D-based experiments, input images are full-resolution $1600\times 900$. All decoder layers have an embedding dimension of 256 and a feed-forward dimension of 512. The query embeddings and their corresponding query position encodings of the detection queries are randomly initialized, and the initial reference points are estimated from their initial position encoding using a linear layer. To better compare with other works, we use $N_{D}=300$ detection queries and freeze the weights of the image backbone and the FPN during training, following STAR-Track [36].
For all PETR-based experiments, image resolutions are $1600\times 640$. The embedding dimension is 256 and the feed-forward dimension is 2048. In contrast to [7], PETR [8] generates query position encodings using the uniformly initialized reference points while initializing detection queries. In the query propagation phase, we adhere to this setting to update the query position encodings using the updated reference point positions at each timestamp. We found this design choice to be essential for ensuring the effectiveness of the PETR-based trackers. Following PF-Track [37], we use $N_{D}=500$ detection queries and only freeze the image backbone for all PETR experiments during training.
In the original version [20], weights of the FPN in the pre-trained detector’s checkpoint were not correctly loaded when training the end-to-end tracker, which we fixed in this version. Another change compared to [20] is the update of track queries that spawn new tracks. Originally [20], the feature representations of these queries remain unchanged to the initialized features, whereas features of all track queries should be updated after being processed by the transformer decoder. In this refined version, features of all track queries are updated before being propagated into the next frame.
5.2 Dataset and Metrics
@@
DETR3D [7] TBA-Baseline 0.341 1.420 0.477 0.310 466 14680 41240 60191 59.73 M 1024 G 296 ms
TBD-Baseline 0.368 1.390 0.482 0.327 987 13294 40833 60077 63.72 M 1025 G 297 ms
ADA-Track 0.391 1.364 0.524 0.343 842 14790 37923 63132 63.73 M 1054 G 308 ms
PETR [8] TBA-Baseline 0.419 1.329 0.532 0.391 231 13996 37425 64241 82.58 M 2076 G 443 ms
TBD-Baseline 0.473 1.246 0.580 0.423 898 14684 32884 68115 97.00 M 2079 G 445 ms
ADA-Track 0.493 1.208 0.602 0.450 672 14647 30422 70803 97.22 M 2140 G 453 ms
Table I: Comparison of ADA-Track with TBA- and TBD-baselines on the nuScenes validation set, grouped by DETR3D and PETR detectors. The inference time per frame (Inf. time) is measured on an RTX 2080ti GPU.
[…]
MUTR3D [35] R101 DETR3D TBA 0.294 1.498 0.427 0.267 3822
MUTR3D [35] R101 DETR3D TBA 0.341 1.420 0.477 0.310 466 14680 41240 60191
DQTrack [17] R101 DETR3D TBD 0.367 1.351 1120
STAR-Track [36] R101 DETR3D TBA 0.379 1.358 0.501 0.360 372
ADA-Track§ [20] R101 DETR3D ADA 0.378 1.391 0.507 0.343 981 15443 38466 62450
ADA-Track R101 DETR3D ADA 0.391 1.364 0.524 0.343 842 14790 37923 63132
ADA-Track++ R101 DETR3D ADA 0.398 1.336 0.544 0.357 774 15345 36381 64742
ADA-Track V2-99 DETR3D ADA 0.423 1.298 0.550 0.391 732 12543 35869 65296
ADA-Track++ V2-99 DETR3D ADA 0.434 1.304 0.556 0.386 863 14723 34529 66505
ADA-Track ViT-L DETR3D ADA 0.512 1.190 0.637 0.462 793 15228 29280 71824
ADA-Track++ ViT-L DETR3D ADA 0.522 1.178 0.630 0.465 664 15444 29275 71958
MUTR3D [35] V2-99 PETR TBA 0.407 1.357 0.511 0.369 271 14401 39487 62139
PF-Track [37] V2-99 PETR TBA 0.479 1.227 0.590 0.435 181
ADA-Track§ [20] V2-99 PETR ADA 0.479 1.246 0.602 0.430 767 15385 31402 69728
ADA-Track V2-99 PETR ADA 0.493 1.208 0.602 0.450 672 14647 30422 70803
ADA-Track++ V2-99 PETR ADA 0.504 1.197 0.608 0.445 613 14839 30616 70668
Table II: Comparison of ADA-Track with existing works on the nuScenes validation set. denotes reproduced version. § denotes previous version [20]. denotes total training epochs of 48. indicates annotations from future frames are required.
@@
ADA-Track++ 0.500 1.144 0.595 0.456 658
Table III: Results on the nuScenes test split. The bottom part compares our method with other end-to-end query-based methods that use DETR3D or PETR. The top part shows other methods.
@@ 5.3.1 Baselines
5.3.2 Results
Table I compares ADA-Track with TBA- and TBD-Baselines on the nuScenes validation split. For both detectors, the TBD-Baselines achieve significantly higher AMOTA than the TBA-Baselines. On top, ADA-Track outperforms the TBD-Baselines by 2.3%P AMOTA for DETR3D and 2.0%P for PETR, respectively, while achieving a considerably higher recall. Considering the secondary metrics, we observe that the TBA-Baseline produces fewer IDS than the TBD-Baseline and ADA-Track, i.e., both methods with the explicit association. However, TBA-Baseline achieves a low IDS by tracking only the easy cases, resulting in a substantially higher FN and lower MOTA. Compared to TBD-Baseline, ADA-Track again improves MOTA by a much higher TP and lower FN. These observations show the importance of the decoupled queries with explicit association as in the TBD-Baseline and in ADA-Track, which produces distinguishable query representations for both tasks and thus improves performances. ADA-Track further improves over the TBD-Baseline by fully utilizing the task inter-dependency using alternating detection and association.
We compare the number of parameters, FLOPs, and runtime of ADA-Track with TBA-Baseline and TBD-Baseline in the right part of Table I. Although ADA-Track and TBD-Baseline require additional edge-augmented cross-attention modules compared to TBA-Baseline, for DETR3D, ADA-Track only adds about $6.7\%$ parameters, $2.9\%$ FLOPs, and $4.1\%$ inference time, and even less compared to TBD. The same tendency can be observed in PETR-based experiments. Despite this slight increase in complexity, the performance gain of ADA-Track is much more significant, e.g., $14.6\%$ and $17.7\%$ higher AMOTA than TBA-Baseline with DETR3D and PETR, respectively.
@@ 5.4 Comparison with Existing Works
We compare ADA-Track and ADA-Track++ with existing end-to-end tracking methods in Table II. First of all, our refined implementation (depicted as ADA-Track) achieved significantly better performance than reported in the preliminary version [20] (depicted as ADA-Track§) for both DETR3D and PETR, especially higher AMOTA and lower IDS. This is due to the different checkpoints of the pre-trained detectors as described in Section 5.1. For the DETR3D-based experiments, MUTR3D [35] was already used as TBA-Baseline in Section 5.3. Interestingly, the original work reported a lower performance than our reproduction with fewer training epochs and a fixed backbone during training. DQ-Track [17] also uses decoupled queries and a sophisticated learned association module following TBD. ADA-Track outperforms it by 2.4%P AMOTA, which again highlights the superiority of our alternating detection and association design. The training of STAR-Track [36] requires an initialization on a pre-trained MUTR3D checkpoint, resulting in a total training epoch of 48. Our ADA-Track can still outperform it by $1.2\%$P AMOTA and achieves a much higher recall. Using larger backbones, i.e., VoVNet-99 and ViT-Large, yields considerable AMOTA increases, demonstrating the generalization to different backbones of ADA-Track. We further observe an overall performance gain of ADA-Track++ compared to ADA-Track for all backbones, highlighting the effectiveness of introducing the auxiliary token in data association. For the PETR-based methods, PF-Track [37] is a joint tracking and prediction method, utilizing a track extension module to replace low-confidence detections with predicted trajectories when outputting tracking results, which additionally requires supervision from future frames. Without any supervision signals from the future, ADA-Track is a pure tracking method but outperforms PF-Track by $1.4\%P$ AMOTA. ADA-Track++ improves ADA-Track by $1.1\%P$ in AMOTA, demonstrating the consistent benefit of the auxiliary token in ADA-Track++.
@@
2 0.321 1.429 0.478 0.286 1696 1.22 s 4.04 GB
3 0.391 1.364 0.524 0.343 842 1.44 s 7.07 GB
4 0.394 1.357 0.524 0.358 766 2.15 s 10.08 GB
Table IV: Ablation study on the training sample length $T$. Time indicates the training time for each step. Memory indicates the memory demand during training.
@@
6 1 0.351 1.390 0.476 0.305 1749
6 2 0.378 1.371 0.494 0.329 1205
6 3 0.389 1.361 0.516 0.344 1051
6 4 0.389 1.359 0.519 0.340 888
6 5 0.390 1.358 0.522 0.344 820
1 6 0.337 1.422 0.489 0.305 965
2 6 0.379 1.373 0.491 0.336 942
3 6 0.389 1.363 0.516 0.343 925
4 6 0.390 1.365 0.509 0.343 847
5 6 0.387 1.368 0.502 0.343 868
6 6 0.391 1.364 0.524 0.343 842
Table V: Ablation study on combining bounding box and association outputs from different decoder layers.
5.5 Ablation Study on ADA-Track
In this section, we report several ablation studies on ADA-Track with a focus on understanding the proposed alternating detection and association paradigm.
5.5.1 Number of training frames
Table IV shows the impact of varying the training sample length $T$. We observe that an increasing sample length increases the AMOTA and many secondary metrics. A particularly substantial improvement occurs when increasing from a sample length of 2 to 3 frames, resulting in a notable 7.0%P increase in AMOTA. The reason is that the autoregressive training scheme becomes active when $T\geq 3$, allowing association results to propagate into subsequent frames. This enables optimization through gradients across multiple frames to optimize the whole sequence, significantly boosting robustness during inference. A further increase in training frames from 3 to 4 yields marginal improvements but a considerable computational and memory overhead during training. Thus, we use $T=3$ as the default setting.
5.5.2 Joint optimization
One of the main contributions of this work is the introduction of an alternating detection and association paradigm, where both tasks iteratively inform each other layer-by-layer for joint optimization. To assess its effectiveness, we combine the predictions of the bounding boxes and the association scores from different decoder layers. The first part of Table V illustrates using bounding boxes from the last layer alongside association scores from decoder layers 1 to 5. A notable increase in AMOTA is observed when using association scores from the second instead of the first layer. Using association scores from higher layers leads to gradually increased AMOTA. We also observe a considerable reduction in IDS. Together, this shows the iterative improvements of the association. The second part of Table V presents results using boxes from various layers combined with association scores from the last layer. We can observe a similar tendency in AMOTA increase with higher layers as in the first part, where the most significant increase occurs from the first to the second layer. In addition, using box predictions in later layers results in a notable improvement in AMOTP, confirming the iterative optimization of localization precision. Overall, combining outputs from distinct layers yields gradual performance increases when outputs from higher layers are used. This observation shows that the relevance between box and association predictions is not only constrained within the same layer but across different layers, which validates the iterative optimization through stacked layers of both tasks.
@@ 5.5.2 Joint optimization
0.0 0.391 1.364 0.524 0.343 842
0.3 0.387 1.363 0.520 0.347 725
0.5 0.384 1.365 0.503 0.341 708
0.7 0.387 1.357 0.520 0.358 793
1.0 0.382 1.369 0.502 0.341 696
Table VI: Impact of the feature update weight $w_{t}$.
@@
A 0.387 1.357 0.510
B 0.378 1.347 0.483
C 0.391 1.364 0.524
Table VII: Ablation study on box update. use track box denotes that boxes predicted from track queries are used as output. track as query denotes that track queries are regarded as queries and detection queries as keys in the edge-augmented cross-attention.
5.5.3 Track update